# Orbit-based structural decomposition and stellar population recovery for edge-on barred galaxies

### v1

@@
Yunpeng Jin E-mail: jyp199333@163.com11    Ling Zhu E-mail: lzhu@shao.ac.cn22    Behzad Tahmasebzadeh 3344    Shude Mao 11    Glenn van de Ven 55    Timothy A. Davis 66
In our previous paper, we developed an orbit-superposition method for edge-on barred galaxies, and constructed dynamical models based on different mock observations of three galaxies from the Auriga simulations. In this study, we adopt 12 cases with side-on bars (three simulated galaxies, each with four different projections). We decompose these galaxies into different structures combining the kinematic and morphological properties of stellar orbits, and compare the model-predicted components to their true counterparts in the simulations. Our models can identify (BP/X-shaped) bars, spheroidal bulges, thin disks, and spatially diffuse stellar halos. The mass fractions of bars and disks are well constrained with absolute biases $|f_{\rm model}-f_{\rm true}|\leq 0.15$. The mass fractions of halos are recovered with $|f_{\rm model}-f_{\rm true}|\leq 0.03$. For the bulge components, 10 of 12 cases have $|f_{\rm model}-f_{\rm true}|\leq 0.05$, while the other two cases have $|f_{\rm model}-f_{\rm true}|\leq 0.10$. Then, by tagging the stellar orbits with ages and metallicities, we derive the chemical properties of each structure. For stellar ages, our models recover the negative gradients in the bars and disks, but exhibit relatively larger uncertainties for age gradients in the bulges and halos. The mean stellar ages of all components are constrained with absolute biases $|t_{\rm model}-t_{\rm true}|\rm\lesssim 1\,Gyr$. For stellar metallicities, our models reproduce the steep negative gradients of the bars and bulges, and all different kinds of metallicity gradients in the disks and halos. Except for the bulge in the simulated galaxy Au-18, the mean stellar metallicities of all other components are constrained with absolute biases $|Z_{\rm model}-Z_{\rm true}|\rm\leq 0.5\,Z_{\odot}$.
@@ 1 Introduction
Our preliminary understanding of galaxy structures comes from their images. A typical spiral galaxy consists of a disk, a bar, and/or a bulge. Since the 1970s, different single-band surveys from ultraviolet to far-infrared wavelengths have been carried out, which provide information about the colour and star formation of these structures. These surveys have indicated that, in spiral galaxies, bulges are typically red, gas-poor, and lack ongoing star formation, while disks are usually blue, gas-rich, and contain different star-forming regions (Kennicutt, 1998). Furthermore, the presence of a bar may enhance star formation in the central region of a galaxy (e.g. Ho et al., 1997a, b; Sakamoto et al., 1999). These findings demonstrated the importance of analysing distinct structures independently in order to better understand galaxy formation and evolution.
@@
In the past decades, significant progress has been made in decomposing galaxy structures. The most common approach, photometric decomposition, relies on the assumption that different components differ in their light distributions. Photometric decomposition was initially performed by fitting 1D surface brightness profiles, considering only two structural components: the bulge and the disk. The bulge component was modelled using the $R^{1/4}$ law (de Vaucouleurs, 1948) or a more generalised Sérsic profile $R^{1/n}$ with $n\gtrsim 2$ (Sersic, 1968), while the disk component was characterised by an exponential profile (e.g. Freeman, 1970; Kormendy, 1977; Burstein, 1979; Boroson, 1981; Kent, 1985). Since the 1990s, high-resolution telescopes such as the Hubble Space Telescope and large-scale sky surveys such as the Sloan Digital Sky Survey (SDSS; York et al., 2000) have provided galaxy images across various wavelengths. These images greatly advanced the development and application of 2D photometric decomposition techniques (e.g. Andredakis et al., 1995; de Jong, 1996; Simard, 1998; Simard et al., 2002; Ratnatunga et al., 1999; Peng et al., 2002, 2010; de Souza et al., 2004; Laurikainen et al., 2004, 2005; Gadotti, 2008; Méndez-Abreu et al., 2008, 2017; Kruk et al., 2018), where the central bulge and bar are either considered as a single component or modelled separately. However, the results of photometric decomposition significantly depend on the assumed analytical functions of the light distributions, which might diverge from the truth. For example, disks are usually modelled with exponential profiles, whereas recent studies have demonstrated that disks can be down-bending or up-bending in their inner or outer regions (Méndez-Abreu et al., 2017; Breda et al., 2020; Ding et al., 2023); bars are often assumed to be prolate Ferrers bars (see Binney & Tremaine, 2008) or are characterised by Sérsic profiles, while real bars can be triaxial with boxy, peanut, or X-shaped (hereafter BP/X-shaped) structures (e.g. Shaw, 1987; Lütticke et al., 2000; Erwin & Debattista, 2017; Li et al., 2017). Furthermore, photometric decomposition provides limited information on the stellar populations of different components.
@@
With the development of integral field spectroscopy (IFS) surveys in recent years, other decomposition techniques have become possible. IFS surveys such as SAURON (Bacon et al., 2001), $\rm ATLAS^{3D}$ (Cappellari et al., 2011), CALIFA (Sánchez et al., 2012), SAMI (Bryant et al., 2015), MaNGA (Bundy et al., 2015), and multiple VLT/MUSE programmes (Bacon et al., 2017) have provided spatially resolved spectra for thousands of nearby galaxies. Therefore, one choice is to decompose galaxies directly by fitting their spectra (e.g. Johnston et al., 2012, 2014; Tabor et al., 2017, 2019; Oh et al., 2020; Pak et al., 2021) based on the full spectral fitting technique such as pPXF (Cappellari & Emsellem, 2004; Cappellari, 2017). This spectroscopic decomposition method can not only separate different structures but also estimate the stellar population of each component. Another approach is to decompose galaxies by combining the photometry and kinematics derived from spectra (e.g. Rigamonti et al., 2023, 2024). However, bar structures are not treated as individual components in these works.
@@
Dynamical decomposition is an alternative technique. This technique is widely used in simulations as the 6D phase-space information of stellar particles is known (e.g. Correa et al., 2017; Pillepich et al., 2019; Rodriguez-Gomez et al., 2019; Du et al., 2019, 2020; Pulsoni et al., 2020). For real galaxies, dynamical decomposition can be achieved through the Schwarzschild’s orbit-superposition method (Schwarzschild, 1979, 1982, 1993), which builds 3D galaxy models by fitting the luminosity distributions and stellar kinematics derived from IFS data. Several implementations of this method are currently in use (e.g. van den Bosch et al., 2008; Long & Mao, 2018; Vasiliev & Valluri, 2020; Neureiter et al., 2021; Quenneville et al., 2022). The van den Bosch et al. (2008) implementation and its publicly available version DYNAMITE (Jethwa et al., 2020) [^1] https://dynamics.univie.ac.at/dynamite_docs/ were validated with simulated galaxies (e.g. Zhu et al., 2018c; Jin et al., 2019) and were used to analyse orbital structures for real galaxies in various IFS surveys (e.g. CALIFA, Zhu et al., 2018a, b, c; MaNGA, Jin et al., 2020; SAMI, Santucci et al., 2022, 2023; $\rm ATLAS^{3D}$, Thater et al., 2023). This implementation was further updated to a population-orbit superposition method, which supports the modelling of stellar populations by fitting the ages and metallicities derived from IFS spectra (e.g. Poci et al., 2019, 2021; Zhu et al., 2020, 2022a, 2022b; Ding et al., 2023; Jin et al., 2024).
@@
Based on the van den Bosch et al. (2008) implementation, a barred orbit-superposition method was developed (Tahmasebzadeh et al., 2021, 2022; see also the public code DYNAMITE), and was applied to a barred galaxy NGC 4371 observed at an inclination angle $\theta\sim 60^{\circ}$ (Tahmasebzadeh et al., 2024). This method is limited to non-edge-on cases ($\theta\lesssim 80^{\circ}$) and stellar populations are not considered. A recent VLT/MUSE programme, GECKOS (van de Sande et al., 2024), is designed to observe 36 Milky Way-like edge-on ($\theta\gtrsim 85^{\circ}$) galaxies and will provide high spatial resolution ($\rm<0.2\,kpc$) IFS data of these galaxies, including both stellar kinematics and stellar populations. Eight of the 12 galaxies in the first GECKOS internal data release are barred and have BP/X-shaped structures (Fraser-McKelvie et al., 2025). Edge-on observations preserve more kinematic details of bar and disk rotations than non-edge-on cases, making them more effective for dynamical decomposition. In our previous paper (Jin et al., 2025), we extended the method developed by Tahmasebzadeh et al. (2021, 2022) to model edge-on ($\theta\gtrsim 80^{\circ}$) barred galaxies that may exhibit BP/X-shaped structures. Using three simulated galaxies from the Auriga simulations (Grand et al., 2017, 2024) and creating a total of 16 mock data sets, the effectiveness of this method in estimating the bar pattern speed $\rm\Omega_{p}$ was validated, achieving an accuracy of $\sim 13\%$. In this paper, we perform orbit-based structural decomposition for edge-on barred galaxies based on the model results from Jin et al. (2025). Then, by tagging the stellar orbits with ages and metallicities, we evaluate how accurately ages and metallicities can be recovered for different components.
@@
The paper is organised as follows. In Sect. 2, we introduce the simulated galaxies and subsequent mock data. In Sect. 3, we demonstrate how we create barred population-orbit superposition models, including the orbit-based decomposition of galaxy structures. In Sects. 4 and 5, we characterise the properties of model-decomposed structures and compare them with their true counterparts in simulations. We discuss our results in Sect. 6 and summarise in Sect. 7.
@@ 2 Auriga simulated galaxies and mock data
Figure 1: Mock stellar kinematic, age, and metallicity maps (left panels), and their corresponding error maps (right panels) for Au-23 with viewing angles $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ})$ with $\rm 1\,arcsec=0.2\,kpc$. From top to bottom: Mean velocity $V$, velocity dispersion $\sigma$, third-order $h_{3}$ and fourth-order $h_{4}$ Gauss-Hermite coefficients, stellar ages $t$, and stellar metallicity $Z$. The kinematic maps and their error maps are the same as Fig. 2 of Jin et al. (2025).
@@
The Auriga project is a suite of cosmological magnetohydrodynamical zoom-in simulations that consists of 30 isolated Milky Way-mass halos (Grand et al., 2017), which are selected from the dark matter-only version of the EAGLE Ref-L100N1504 simulation (Schaye et al., 2015). We adopt three simulated galaxies from the Auriga simulations: Au-18, Au-23, and Au-28, which are the same as in our previous paper (Jin et al., 2025). Au-18 and Au-23 have clear BP/X-shaped structures, while Au-28 does not. The stellar masses of these three galaxies are $M_{*}=8.04\times 10^{10}\,\rm M_{\odot}$, $9.02\times 10^{10}\,\rm M_{\odot}$, and $10.45\times 10^{10}\,\rm M_{\odot}$, respectively, and their corresponding dark matter halo masses are $M_{\rm 200}=1.22\times 10^{12}\,\rm M_{\odot}$, $1.58\times 10^{12}\,\rm M_{\odot}$, and $1.61\times 10^{12}\,\rm M_{\odot}$.
@@
For the three simulated galaxies, we adopt the mock surface brightness and stellar kinematic maps created in Jin et al. (2025). Each simulated galaxy is observed with four different projections $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ})$, $(85^{\circ},80^{\circ})$, $(89^{\circ},50^{\circ})$, and $(89^{\circ},80^{\circ})$, where $\theta_{\rm T}$ is the inclination angle and $\varphi_{\rm T}$ is the bar azimuthal angle ($\varphi_{\rm T}=0^{\circ}$ represents that the bar is perfectly end-on, while $\varphi_{\rm T}=90^{\circ}$ denotes that the bar is perfectly side-on). The redshift $z=0$ snapshots were adopted and the galaxies were placed at a distance of $\rm 41.25\,Mpc$ ($\rm 1\,arcsec=0.2\,kpc$), which is similar to the median distance of the GECKOS sample ($15<D\,\rm[Mpc]<70$, van de Sande et al., 2024). The mock surface brightness was constructed from the projected stellar mass distribution by assuming a constant stellar mass-to-light ratio $M_{*}/L=2$ and a pixel size of $\rm 1\times 1\,arcsec^{2}$. The pixels were then binned to MUSE-like apertures using the Voronoi binning method (Cappellari & Copin, 2003). The mock kinematic maps were derived by fitting the velocity distribution of the stellar particles in each aperture with a Gauss-Hermite expansion (Gerhard, 1993; van der Marel & Franx, 1993; Rix et al., 1997). We refer to Jin et al. (2025) for the details of creating these mock data. We have 12 mock data sets (three simulated galaxies, each with four different projections) in total. Each mock data set is labelled with the format ‘name-$\theta_{\rm T}$-$\varphi_{\rm T}$’, for example, the mock observation of Au-23 with viewing angles $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ})$ is denoted as Au-23-85-50. This notation is used consistently in the paper.
@@
Here we further create mock stellar age maps and stellar metallicity maps for each projection $(\theta_{\rm T},\varphi_{\rm T})$ of each galaxy. For both age and metallicity maps, we use the same apertures as the kinematic maps. Within each aperture $i$, we determine the mean age $t_{i}$ and the mean metallicity $Z_{i}/Z_{\odot}$ by calculating the mass-weighted values of all stellar particles located within the aperture. We assume the percentage errors of the age and metallicity maps to be $10\%$, and add Gaussian perturbations to the maps: $t_{i}^{\prime}=t_{i}+0.1t_{i}a$, and $Z_{i}^{\prime}=Z_{i}+0.1Z_{i}b$, where $a$ and $b$ are random values with standard deviations equal to unity. The perturbed age maps $t_{i}^{\prime}$ and metallicity maps $Z_{i}^{\prime}/Z_{\odot}$ are taken as mock observations, while $0.1t_{i}$ and $0.1Z_{i}$ are treated as corresponding error maps. Fig. 1 illustrates the stellar kinematics, ages, and metallicities for Au-23-85-50.

### v4

@@
Yunpeng Jin E-mail: jyp199333@163.com    Ling Zhu E-mail: lzhu@shao.ac.cn    Behzad Tahmasebzadeh    Shude Mao    Glenn van de Ven    Timothy A. Davis
In our previous paper, we developed an orbit-superposition method for edge-on barred galaxies and constructed a set of dynamical models based on different mock observations of three galaxies from the Auriga simulations. In this study, we adopted 12 cases with side-on bars (three simulated galaxies, each with four different projections). We decomposed these galaxies into different structures combining the kinematic and morphological properties of stellar orbits. We then compared the model-predicted components to their true counterparts in the simulations. Our models can identify (BP/X-shaped) bars, spheroidal bulges, thin discs, and spatially diffuse stellar halos. The mass fractions of bars and discs are well constrained with absolute biases of $|f_{\rm model}-f_{\rm true}|\leq 0.15$. We recovered the mass fractions of halos with $|f_{\rm model}-f_{\rm true}|\leq 0.03$. For the bulge components, 10 out of 12 cases exhibit $|f_{\rm model}-f_{\rm true}|\leq 0.05$, while the other two cases exhibit $|f_{\rm model}-f_{\rm true}|\leq 0.10$. Then, by tagging the stellar orbits with ages and metallicities, we derived the chemical properties of each structure. For the stellar ages, our models recovered the negative gradients in the bars and discs, but exhibited relatively larger uncertainties for age gradients in the bulges and halos. The mean stellar ages of all components were constrained with absolute biases $|t_{\rm model}-t_{\rm true}|\rm\lesssim 1\,Gyr$. For stellar metallicities, our models reproduced the steep negative gradients of the bars and bulges, as well as all different kinds of metallicity gradients in the discs and halos. Apart from the bulge in the simulated galaxy Au-18, the mean stellar metallicities of all other components were constrained with absolute biases of $|Z_{\rm model}-Z_{\rm true}|\rm\leq 0.5\,Z_{\odot}$.
@@ 1 Introduction
Our preliminary understanding of galaxy structures comes from their images. A typical spiral galaxy consists of a disc, a bar, and/or a bulge. Since the 1970s, different single-band surveys from ultraviolet (UV) to far-infrared (FIR) wavelengths have been carried out, providing information about the colour and star formation of these structures. These surveys have indicated that in spiral galaxies, bulges are typically red, gas-poor, and lack ongoing star formation, while discs are usually blue, gas-rich, and contain different star-forming regions (Kennicutt, 1998). Furthermore, the presence of a bar may enhance star formation in the central region of a galaxy (e.g. Ho et al., 1997a, b; Sakamoto et al., 1999). These findings demonstrated the importance of analysing distinct structures independently to better understand galaxy formation and evolution.
@@
In the past decades, significant progress has been made in decomposing galaxy structures. The most common approach, photometric decomposition, relies on the assumption that different components differ in their light distributions. Photometric decomposition was initially performed by fitting 1D surface brightness profiles, considering only two structural components: the bulge and the disc. The bulge component was modelled using the $R^{1/4}$ law (de Vaucouleurs, 1948) or a more generalised Sérsic profile, $R^{1/n}$, with $n\gtrsim 2$ (Sersic, 1968), while the disc component was characterised by an exponential profile (e.g. Freeman, 1970; Kormendy, 1977; Burstein, 1979; Boroson, 1981; Kent, 1985). Since the 1990s, high-resolution telescopes such as the Hubble Space Telescope and large-scale sky surveys such as the Sloan Digital Sky Survey (SDSS; York et al., 2000) have provided galaxy images across various wavelengths. These images greatly advanced the development and application of 2D photometric decomposition techniques (e.g. Andredakis et al., 1995; de Jong, 1996; Simard, 1998; Simard et al., 2002; Ratnatunga et al., 1999; Peng et al., 2002, 2010; de Souza et al., 2004; Laurikainen et al., 2004, 2005; Gadotti, 2008; Méndez-Abreu et al., 2008, 2017; Kruk et al., 2018), where the central bulge and bar are either considered as a single component or modelled separately. However, the results of photometric decomposition significantly depend on the assumed analytical functions of the light distributions, which might diverge from the truth. For example, discs are usually modelled with exponential profiles, whereas recent studies have demonstrated that discs can be down-bending and/or up-bending in their inner or outer regions (Méndez-Abreu et al., 2017; Breda et al., 2020; Ding et al., 2023); bars are often assumed to be prolate Ferrers bars (see Binney & Tremaine, 2008) or are characterised by Sérsic profiles, while real bars can be triaxial with boxy, peanut, or X-shaped (hereafter BP/X-shaped) structures (e.g. Shaw, 1987; Lütticke et al., 2000; Erwin & Debattista, 2017; Li et al., 2017). Furthermore, photometric decomposition provides limited information on the stellar populations of different components.
@@
With the development of integral field spectroscopy (IFS) surveys in recent years, other decomposition techniques have become possible. Such IFS surveys as SAURON (Bacon et al., 2001), $\rm ATLAS^{3D}$ (Cappellari et al., 2011), CALIFA (Sánchez et al., 2012), SAMI (Bryant et al., 2015), MaNGA (Bundy et al., 2015), and multiple VLT/MUSE programmes (Bacon et al., 2017) have provided spatially resolved spectra for thousands of nearby galaxies. Therefore, one option is to decompose galaxies directly by fitting their spectra (e.g. Johnston et al., 2012, 2014; Tabor et al., 2017, 2019; Oh et al., 2020; Pak et al., 2021) based on the full spectral fitting technique such as pPXF (Cappellari & Emsellem, 2004; Cappellari, 2017). This spectroscopic decomposition method is not only capable of separating different structures, but also estimating the stellar population of each component. Another approach is to decompose galaxies by combining the photometry and kinematics derived from the spectra (e.g. Rigamonti et al., 2023, 2024). However, bar structures are not treated as individual components in these works.
@@
Dynamical decomposition is an alternative technique. This technique is widely used in simulations as the 6D phase-space information of stellar particles is known (e.g. Correa et al., 2017; Pillepich et al., 2019; Rodriguez-Gomez et al., 2019; Du et al., 2019, 2020; Pulsoni et al., 2020). For real galaxies, dynamical decomposition can be achieved through the Schwarzschild’s orbit-superposition method (Schwarzschild, 1979, 1982, 1993), which builds 3D galaxy models by fitting the luminosity distributions and stellar kinematics derived from IFS data. Several implementations of this method are currently in use (e.g. van den Bosch et al., 2008; Long & Mao, 2018; Vasiliev & Valluri, 2020; Neureiter et al., 2021; Quenneville et al., 2022; Dattathri et al., 2024). The van den Bosch et al. (2008) implementation and its publicly available version DYNAMITE (Jethwa et al., 2020) [^1] https://dynamics.univie.ac.at/dynamite_docs/ were validated with simulated galaxies (e.g. Zhu et al., 2018c; Jin et al., 2019) and were used to analyse orbital structures for real galaxies in various IFS surveys (e.g. CALIFA, Zhu et al., 2018a, b, c; MaNGA, Jin et al., 2020; SAMI, Santucci et al., 2022, 2023; $\rm ATLAS^{3D}$, Thater et al., 2023). This implementation was further updated to a population-orbit superposition method, which supports the modelling of stellar populations by fitting the ages and metallicities derived from IFS spectra (e.g. Poci et al., 2019, 2021; Zhu et al., 2020, 2022a, 2022b; Ding et al., 2023; Jin et al., 2024).
@@
Based on the van den Bosch et al. (2008) implementation, a barred orbit-superposition method was developed (Tahmasebzadeh et al., 2021, 2022; see also the public code DYNAMITE) and applied to a barred galaxy NGC 4371, observed at an inclination angle $\theta\sim 60^{\circ}$ (Tahmasebzadeh et al., 2024). This method is limited to non-edge-on cases ($\theta\lesssim 80^{\circ}$) and stellar populations are not considered. The recently launched VLT/MUSE programme, GECKOS (van de Sande et al., 2024), is designed to observe 36 Milky Way-like edge-on ($\theta\gtrsim 85^{\circ}$) galaxies. It will provide high spatial resolution ($\rm<0.2\,kpc$) IFS data of these galaxies, including both stellar kinematics and stellar populations. Eight of the 12 galaxies in the first GECKOS internal data release are barred and have BP/X-shaped structures (Fraser-McKelvie et al., 2025). Edge-on observations preserve more kinematic details of bar and disc rotations than non-edge-on cases, making them more effective for dynamical decomposition. In our previous paper (Jin et al., 2025), we extended the method developed by Tahmasebzadeh et al. (2021, 2022) to model edge-on ($\theta\gtrsim 80^{\circ}$) barred galaxies that may exhibit BP/X-shaped structures. Using three simulated galaxies from the Auriga simulations (Grand et al., 2017, 2024) and creating a total of 16 mock datasets, the effectiveness of this method in estimating the bar pattern speed, $\rm\Omega_{p}$, was validated, achieving an accuracy of $\sim 13\%$. In this paper, we perform orbit-based structural decomposition for edge-on barred galaxies based on the model results from Jin et al. (2025). Then, by tagging the stellar orbits with ages and metallicities, we evaluate how accurately ages and metallicities can be recovered for different components.
@@
The paper is organised as follows. In Sect. 2, we introduce the simulated galaxies and subsequent mock data. In Sect. 3, we demonstrate how we create barred population-orbit superposition models, including the orbit-based decomposition of galaxy structures. In Sects. 4 and 5, we characterise the properties of model-decomposed structures and compare them with their true counterparts in simulations. We discuss our results in Sect. 6 and summarise our conclusions in Sect. 7.
@@ 2 Auriga simulated galaxies and mock data
Figure 1: Mock stellar kinematic, age, and metallicity maps (left panels), and their corresponding error maps (right panels) for Au-23 with viewing angles, $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ}),$ with $\rm 1\,arcsec=0.2\,kpc$. From top to bottom: Mean velocity, $V$; velocity dispersion, $\sigma$; third-order, $h_{3}$, and fourth-order, $h_{4}$, Gauss-Hermite coefficients; stellar age, $t$; and stellar metallicity, $Z$. The kinematic maps and their error maps are the same as Fig. 2 of Jin et al. (2025). We note that the kinematic error maps exhibit similar statistical properties as they are generated from particle noise (Tsatsi et al., 2015), but their ranges are different (as indicated by the colour bars in the right panels).
@@
The Auriga project is a suite of cosmological magnetohydrodynamical zoom-in simulations that consists of 30 isolated Milky Way-mass halos (Grand et al., 2017), which are selected from the dark matter-only version of the EAGLE Ref-L100N1504 simulation (Schaye et al., 2015). We adopted three simulated galaxies from the Auriga simulations: Au-18, Au-23, and Au-28, which are the same as in our previous paper (Jin et al., 2025). Au-18 and Au-23 have clear BP/X-shaped structures, while Au-28 does not. The stellar masses of these three galaxies are $M_{*}=8.04\times 10^{10}\,\rm M_{\odot}$, $9.02\times 10^{10}\,\rm M_{\odot}$, and $10.45\times 10^{10}\,\rm M_{\odot}$, respectively, and their corresponding dark matter halo masses are $M_{\rm 200}=1.22\times 10^{12}\,\rm M_{\odot}$, $1.58\times 10^{12}\,\rm M_{\odot}$, and $1.61\times 10^{12}\,\rm M_{\odot}$.
@@
For the three simulated galaxies, we adopted the mock surface brightness and stellar kinematic maps created in Jin et al. (2025). Each simulated galaxy was observed with four different projections $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ})$, $(85^{\circ},80^{\circ})$, $(89^{\circ},50^{\circ})$, and $(89^{\circ},80^{\circ})$, where $\theta_{\rm T}$ is the inclination angle and $\varphi_{\rm T}$ is the bar azimuthal angle ($\varphi_{\rm T}=0^{\circ}$ represents that the bar is perfectly end-on, while $\varphi_{\rm T}=90^{\circ}$ denotes that the bar is perfectly side-on). The redshift $z=0$ snapshots were adopted and the galaxies were placed at a distance of $\rm 41.25\,Mpc$ ($\rm 1\,arcsec=0.2\,kpc$), which is similar to the median distance of the GECKOS sample ($15<D\,\rm[Mpc]<70$; van de Sande et al., 2024).
The mock surface brightness was constructed from the projected stellar mass distribution by assuming a constant stellar mass-to-light ratio of $M_{*}/L=2$. We note that different values of the constant $M_{*}/L$ do not affect the results, as luminosity is treated as a scale factor in our models. The surface brightness maps were initially constructed with a pixel size of $\rm 1\times 1\,arcsec^{2}$. These pixels were then adaptively binned using the Voronoi binning method (Cappellari & Copin, 2003) to generate binned apertures that match the typical scale of those in MUSE kinematic maps at $S/N=50$ to 100 (despite our initial pixels being coarser than MUSE’s $\rm 0.2\times 0.2\,arcsec^{2}$ pixels). The mock kinematic maps were derived by fitting the velocity distribution of the stellar particles in each aperture with a Gauss-Hermite expansion (Gerhard, 1993; van der Marel & Franx, 1993; Rix et al., 1997) and were then perturbed by adding Gaussian noise based on particle numbers following the method in Tsatsi et al. (2015). We refer to Jin et al. (2025) for the details of creating these mock data. We have 12 mock datasets (three simulated galaxies, each with four different projections) in total. Each mock dataset is labelled with the format ‘name-$\theta_{\rm T}$-$\varphi_{\rm T}$’; for example, the mock observation of Au-23 with viewing angles $(\theta_{\rm T},\varphi_{\rm T})=(85^{\circ},50^{\circ})$ is denoted as Au-23-85-50. This notation is used consistently throughout the paper.
@@
We further created mock stellar age maps and stellar metallicity maps for each projection $(\theta_{\rm T},\varphi_{\rm T})$ of each galaxy. For both the age and metallicity maps, we used the same apertures as the kinematic maps. Within each aperture, $i$, we determined the mean age, $t_{i}$, and the mean metallicity, $Z_{i}/Z_{\odot}$, by calculating the mass-weighted values of all stellar particles located within the aperture. We assumed the percentage errors of the age and metallicity maps to be $10\%$ and added Gaussian perturbations to the maps: $t_{i}^{\prime}=t_{i}+0.1t_{i}a$, and $Z_{i}^{\prime}=Z_{i}+0.1Z_{i}b$, where $a$ and $b$ are random values with standard deviations equal to unity. The perturbed age maps, $t_{i}^{\prime}$, and metallicity maps, $Z_{i}^{\prime}/Z_{\odot}$, were taken as mock observations, while $0.1t_{i}$ and $0.1Z_{i}$ were treated as corresponding error maps. Fig. 1 illustrates the stellar kinematics, ages, and metallicities for Au-23-85-50.